# BihoT: A Large-Scale Dataset and Benchmark for Hyperspectral Camouflaged Object Tracking

### v1

@@ III-C Data Statistics
To further analyze the BihoT, Fig. 4 shows the images’ motion distribution and object size changes. We first calculate the average probability density function map. The brighter area indicates the area where the object centroid frequently appears. As shown in Fig. 4(a), the bright area has a wide range and appears scattered. This indicates that the object motion in our dataset is irregular, not moving uniformly like pedestrians or vehicles on the road, which is more challenging. Then, we calculate the ratio of object pixels to images to calculate the object size. As shown in Fig. 4(b), the BihoT contains objects of different sizes, and small- or medium-sized objects have a higher proportion. On one hand, the size variation of an object is beneficial for testing the ability of the tracker to handle multi-scale objects. On the other hand, small-sized objects are more prone to occlusion and background noise, which poses a higher challenge to the robustness of the tracker. By analyzing the image information, centroid distribution, and object size changes, the characteristics of BihoT can be understood, which motivates the investigation of advanced HCOT methods.
@@ IV Proposed Method
Fig. 5 depicts the framework overview of the proposed SPDAN, which consists of two key components. 1) Spectral Prompt-based Backbone Network: This component designs a spectral embedding module that can help establish rich spectral embedding tokens and learn joint feature representations of HS and visual by embedding spectrum as prompts in the network. Therefore, it can inherit the powerful tracking ability of the visual foundation model while solving the problems of scarcity and poor portability of HS data. 2) Distractor-aware Module: This module solves disturbance interference by first establishing a perceptual disturbance statistic that can capture frames with occlusion or background interference in the video sequence. When the SPBN prediction results are unreliable, a motion model is used to update the object’s position, which can effectively reduce the decrease in the tracking performance caused by disturbances.
@@ IV-A Spectral Prompt-based Backbone Network
2) Cross-modality Adapter (CA). SPBN freezes the entire backbone model and only learns spectral prompts for a specific modality without training the visual Transformer backbone from scratch, enabling the model to converge quickly in the limited sample situations. With the embedding tokens, the spectral prompts can be obtained by a cross-modality adapter. As shown in Fig. 6, the initial prompts can be obtained by and . For the th layer of the Transformer, the prompts learned from the CA can be formally represented as
@@ IV-B Distractor-aware Module
where and denote the maximum and minimum values of the classification map , respectively, and denotes the number of pixels of . This DC reveals the degree of fluctuation in the response map and the confidence level of the detected objects. When this value is large, it indicates that the response map fluctuates, and only one extreme point of the response value. The predicted position is close to the true one. When encountering disturbances caused by changes in conditions, the fluctuation of the response map increases and multiple ambiguous prediction points will appear, causing a decrease in DC. Fig. 7 shows the values of DC and the attention heatmap generated by the Grad-CAM++ [39]. It can be observed that when there are disturbances such as occlusion and similar background interference (frame #0055), the prediction results become unreliable. The attention heatmap shows multiple areas with high responses, resulting in the ambiguous tracking results. In this way, the distractor can be perceived.
@@
where is the covariance matrix, denotes the posterior estimation of state vector, is the identity matrix, denotes the Kalman gain, and denotes the prior estimation of the error matrix. To obtain the robust tracking results further, the Kalman tracker only updates the state predicted by SPBN when the DC exceeds the threshold.
Step 3. Rectifying. During the tracking process, SPBN combined with the Kalman filter can effectively handle the problem of short-term occlusion. However, when the environment or other objects are visually similar to the tracking object, especially when spectral information is similar, the DC is still high, even if there is drift in the tracking box. We call this phenomenon the “Confident Error.” At this point, the tracker chooses to trust the results of SPBN (because of its high DC), but it may lead to incorrect tracking, as shown in Fig. 8. If a perfect tracklet can be obtained, the offset of the predicted box between continuous frames will be small. In practice, there are generally only two situations where a significant offset occurs: one is due to a confident error that causes the tracking box to drift, and the other is due to the re-capture of the object when it reappears after the error result is generated by the distractors. We hope to rectify the situation with a larger offset in the first case and retain the ability of the tracker to capture the object again in the second case. Specifically, before occlusion or distractor occurs, the confidence level often gradually decreases. When confident errors occur, the DC will surge, and the tracker will incorrectly predict the position of the object, causing drift of the prediction box. To address this issue, we propose a tracking rectifying strategy that combines DC and offset. When the box drift occurs, we calculate the trend of DC within the range of the past time sliding window. The confidence level gradually decreases, indicating that confident errors will occur. To handle this situation, we directly use the object state predicted by the Kalman filter as the tracking results. By doing so, we can effectively correct the first situation without the second situation that is correct. Algorithm 1 outlines the whole tracking algorithm.
@@ V-A Experiment Settings
HOTC-2020 Dataset: HOTC-2020 was proposed at the first Hyperspectral Object Tracking Challenge in 2020. It consists of 75 HS video sequences, of which 40 are used for training, and 35 are used for testing. The training part contains a total of 12,839 annotated frames, while the testing part contains 16,574 annotated frames. The dataset is annotated with 11 attributes including Background Clutter (BC), Deformation (DEF), Fast Motion (FM), In-Plane Rotation (IPR), Illumination Variation (IV), Low Resolution (LR), Motion Blur (MB), Occlusion (OCC), Out-of-Plane Rotation (OPR), Out-of-View (OV), and Scale Variation (SV).
2) Evaluation Metrics. We choose the commonly used two evaluation metrics: the area under the curve (AUC) of the success plot and the precision rate at the threshold of 20 pixels to conduct experiments. Success plots are used to evaluate the Intersection over Union (IOU) metric between the predicted and actual bounding boxes. These plots depict the percentage of frames that meet or exceed different IOU thresholds, forming a success curve with thresholds ranging from 0 to 1. The precision rate analyzes the pixel difference between the centers of the predicted and actual bounding boxes. The precision rate curve is generated by displaying the percentage of frames where the center discrepancy is less than a specified pixel threshold, which ranges up to 50 pixels. Typically, the precision plot with the threshold of 20 is used to rank the trackers, denoted as DP_20.
3) Implementation Details. The proposed SPDAN is implemented in PyTorch and trained on a machine equipped with an NVIDIA RTX 4090 Graphics Processing Unit (GPU). The backbone network SPBN and the embedding network of RGB modality are initialized with the pre-trained parameters in [35]. The AdamW optimizer with a weight decay of 0.0001 is adopted to optimize the parameters of the network. The template and search patches are set to pixels and pixels, respectively. The network is then trained for a total of 10 epochs. and the batch size is set to 32. While the DC threshold and offset are set to 0.01 and 20, respectively.
@@
| B+S+D | 0.725 | 0.968 | +0.032 | +0.044 |
@@
V-B Ablation Experiments
1) Effectiveness of the Proposed SEN: The spectral embedding network is proposed to refine the spectral features and form a robust prompt representation. To demonstrate the effectiveness of the SEN, we first construct a prompt-based framework following [16], which is denoted as “B”. Then, the SEN is utilized to incorporate the encoded spectral information into the visual features during the token embedding stage. It is denoted as “S”. We conduct the comparative experiments between the two models on the BihoT and HOTC-2020 datasets. The experimental results are listed in Tables III and IV. It can be observed that the SEN significantly improves the performance compared to the baseline model. Specifically, on the BihoT dataset, the AUC score increases by 3.3%, and the DP_20 pixels score improves by 3.8%. On the HOTC-2020 dataset, the AUC score increases by 1.1%, and the DP_20 pixels score improves by 1.4%. This indicates that spectral information is important for the model to capture camouflaged objects,
2) Effectiveness of the Proposed DAM:
DAM is proposed to perceive the occlusion distractors and rectify the tracking results. We add the DAM to the model to show its effectiveness, which is denoted as “D”. The experimental results are listed in Tables III and IV. It can be observed that DAM further enhances the tracking performance. Specifically, DAM increases the AUC and DP on the BihoT dataset from 0.672 and 0.805 to 0.677 and 0.811, respectively. Compared to the baseline model, there are improvements of 3.8% and 4.4% in both indicators. For the HOTC-2020 dataset, DAM increases the AUC score and DP from 0.704 and 0.938 to 0.725 and 0.968, respectively. Compared with the baseline model, there are improvements of 3.2% and 4.4% in both indicators. These results validate the effectiveness of the proposed DAM, which can correct the tracking results misguided by the distractors through a proposed decision confidence metric and three-step processes.
3) The Impact of Embedding Dimension In the SEN:
To explore the impact of embedding dimension (the number of output channels by 3-D convolutional layer) on the performance of the proposed SEN, we select multiple embedding dimension settings and conduct the tracking performance experiments on the proposed BihoT. The experimental results are listed in Table V. It can be seen that the four sets of embedding parameters (i.e., 2,4,8 and 16) are selected. Among them, the setting of embedding dimension 16 achieves the best results: the AUC score reaches 0.672, and DP_20 reaches 0.805. Under optimal experimental settings, the SEN can capture more comprehensive discriminative information in the spectral dimension, providing richer spectral semantic information for the prompt module.
@@ V-C Comparison with State-of-the-arts
Performance on the BihoT Dataset. We compare the proposed SPDAN with advanced HS and RGB trackers on the BihoT dataset, including SEE-Net [14], SiamHYPER [12], MixFormer [34], SiamBAN [40] and OSTrack [35]. SEE-Net reconstructs the HS bands through an autoencoder to learn the weight of each band. SiamHYPER proposes a Siamese tracking network that adopts independent networks for feature extraction of RGB and HS modalities. MixFormer proposes a mixed attention module to extract features, and the tracking box can be obtained without a certain head network. SiamBAN proposes a box-adaptive head network to predict the accurate object positions. OSTrack is a one-stage tracking framework, which extracts the features of the template and search images simultaneously through a visual Transformer-based backbone. The experimental results are listed in Table VI. It can be observed that the proposed SPDAN achieves the best performance in both metrics. Furthermore, the BihoT dataset contains a large number of camouflage objects, which exhibit the same appearances but different spectral information. The lack of visual features poses a significant challenge to traditional RGB trackers that rely on visual features such as color and texture. This unique setting causes substantial disparities in AUC and DP scores between the conventional RGB trackers and the proposed SPDAN. This discrepancy not only highlights the limitations of current trackers but also emphasizes the effectiveness of our methods. In addition, in the comparison methods, most HS trackers use band regrouping to generate three-channel false-color images and directly use RGB trackers to predict object positions without considering spectral information. They heavily rely on the performance of RGB trackers and are not conducive to improving the performance of HS tracking. However, SiamHYPER achieves commendable results. Unlike band regrouping-based methods, it extracts spectral features and fuses cross-modality features, which, to some extent, utilizes spectral information and alleviates the dependence on RGB trackers. It indicates the importance of spectral information for HOT, especially for camouflaged objects.
@@
Performance on the HOTC-2020 Dataset. In this sub-section, we compare the proposed SPDAN with recent HS and RGB trackers in this sub-section. HS trackers include BAE-Net [45], SST-Net [46], SiamHYPER [12], SEE-Net [14], SiamOHOT [13], SPIRIT[11], MMF-Net [28], and SENSE [10], which are tested on HS video sequences. The RGB trackers include SiamBAN [40], SiamGAT [41], SiamCAR [42], SimTrack [43], SBT [38], and SeqTrack [44], which test on false-color video sequences. The experimental results are shown in Table VII and Figs. 10(a) and (b). Our proposed SPDAN can achieve results of 0.725 on AUC and 0.968 on DP_20, surpassing the above recent methods and proving the effectiveness of our method. The AUC of RGB trackers is mostly below 0.600, and the DP_20 is below 0.900. The performance of HS trackers is generally better than that of RGB trackers. This result proves the necessity of leveraging spectral features, as relying solely on visual features cannot handle distractors from similar objects and background clutter. Then, comparing with HS trackers, it can be observed that the use of cross-modality feature fusion performs better, indicating that fusing RGB and HS modalities can help improve performance. In addition, most existing HS trackers are based on the Siamese architecture of SiamFC [47], ignoring the advantages of the Transformer in the feature extraction. Their global relationship modeling ability is very helpful for extracting the spatial and spectral information. However, ViT-based trackers like SimTrack and SeqTrack face limitations due to their heavy reliance on the visual features and a semantic discrepancy with the robust HS features, impeding their performance enhancement. Their AUC yields 0.600 and 0.594, while the DP_20 are 0.845 and 0.856, respectively. In comparison, our SPDAN has achieved improvements of 12.5% and 13.1% in AUC and 12.3% and 11.2% in DP_20, respectively.
@@
To demonstrate the superiority of our method, we show the tracking results on the proposed BihoT dataset. We select four video sequences, including lemon2, apple3, chill2, and tank1. Due to the lack of RGB images collected in our dataset, we select the 1st, 9th, and 16th bands to form the false-color image. The experimental results are shown in Fig. 9. The ground truth bounding box is framed by black lines. The first column is obtained from the second frame of tracking results. The results from the consecutive frames are illustrated from the second to the fourth columns. From the first and second video sequences, we can observe that most methods struggle to accurately distinguish between camouflaged and real objects, resulting in tracking failures. In contrast, our method successfully captures both visual and spectral information of the objects, demonstrating the robustness in handling camouflaged objects. In addition, for the third video sequence, SPDAN can still perform well for smaller objects, especially when dealing with object rotation. Furthermore, The fourth sequence introduces two types of challenges related to camouflage: the first is the environmental similarity to the tank model, and the second is the visual similarity between two tanks made from different materials. It can be seen that the proposed method can effectively address these two challenges. In comparison, methods like SiamHYPER and MixFormer are misled by the camouflaged tanks during object movement. Other methods are also impacted by background clutter or camouflaged elements, further underscoring the superior performance of our method.
V-E Running Time Comparison.
To demonstrate the superiority of our method in terms of running speed, we conduct experiments on the BihoT dataset. The experimental results using frames per second (FPS) as the evaluation indicator are reported in Table VIII. All the trackers ran on a Windows machine with the configuration of an Intel(R) Core(TM) i7-13700K CPU@3.40GHz and an NVIDIA 4090 GPU. Besides, the RGB trackers ran on the three-channel false-color images, and the HS tracker ran on the HSIs. SEE-Net, SiamHYPER, and SPDAN are HS trackers, and others are RGB trackers. SPDAN achieves the best running speed among all trackers. It proves that our method has better computational efficiency compared to band regrouping or existing modality fusion methods

### v2

@@ IV Proposed Method
Fig. 3 depicts the framework overview of the proposed SPDAN, which consists of two key components. 1) Spectral Prompt-based Backbone Network: This component designs a spectral embedding module that can help establish rich spectral embedding tokens and learn joint feature representations of HS and visual by embedding spectrum as prompts in the network. Therefore, it can inherit the powerful tracking ability of the visual foundation model while solving the problems of scarcity and poor portability of HS data. 2) Distractor-aware Module: This module solves disturbance interference by first establishing a perceptual disturbance statistic that can capture frames with occlusion or background interference in the video sequence. When the SPBN prediction results are unreliable, a motion model is used to update the object’s position, which can effectively reduce the decrease in tracking performance caused by disturbances.
@@ IV-A Spectral Prompt-based Backbone Network
2) Cross-modality Adapter (CA). SPBN freezes the entire backbone model and only learns spectral prompts for a specific modality without training the visual Transformer backbone from scratch, enabling the model to converge quickly in the limited sample situations. A cross-modality adapter (CA) is employed to learn the hybrid prompt tokens from HS and visual modalities, which leverages the fovea attention to extract HS features. This integration allows the model to capture complementary information from both HS and visual modalities, enhancing the ability to distinguish fine-grained spectral details while maintaining spatial coherence. As shown in Fig. 4, the initial prompts can be obtained by and . For the th layer of the Transformer, the prompts learned from the CA can be formally represented as
@@
To effectively train the SPDAN, we adopt the weighted focal loss [47] for classification, and IoU loss [48] and loss for regression
where and are set to 2 and 5, respectively, according to [49].
@@ IV-B Distractor-aware Module
where and denote the maximum and minimum values of the classification map , respectively, and denotes the number of pixels of . This DC reveals the degree of fluctuation in the response map and the confidence level of the detected objects. When this value is large, it indicates that the response map is stable, and there is only one extreme point of the response value. The predicted position is close to the true one. When encountering disturbances caused by changes in conditions, the fluctuation of the response map increases, and multiple ambiguous prediction points will appear, causing a decrease in DC. Fig. 5 shows the values of DC and the attention heatmap generated by the Grad-CAM++ [50]. It can be observed that when there are disturbances such as occlusion and similar background interference (frame #0055), the prediction results become unreliable. The attention heatmap shows multiple areas with high responses, resulting in ambiguous tracking results. In this way, the distractor can be perceived.
@@
where is the covariance matrix, denotes the posterior estimation of state vector, is the identity matrix, denotes the Kalman gain, and denotes the prior estimation of the error matrix. To obtain robust tracking results further, the Kalman tracker only updates the state predicted by SPBN when the DC exceeds the threshold.
Step 3. Rectifying. During the tracking process, SPBN combined with the Kalman filter can effectively handle short-term occlusion. However, an issue arises with the decision confidence (DC) metric proposed in [21]. When the environment or surrounding objects are visually similar to the tracking object, particularly when their spectral information is also similar, the DC remains high, even if the tracking box has drifted. We call this phenomenon a “Confident Error.” In such cases, the tracker trusts the tracing results from SPBN (because of the high DC), leading to incorrect tracking, as shown in Fig. 6. If a perfect tracklet can be obtained, the offset of the predicted box between continuous frames remains small. Significant offsets can arise under two conditions: either a confident error has caused the tracking box to drift, or the object has been successfully re-captured after being lost due to occlusion or other distractors. Our goal is to rectify the situation with a significant offset in the first case while preserving the ability of the tracker to capture the object again in the second case.
Different from [21], we propose a rectifying strategy that combines DC and offset. Specifically, before the distractor occurs, the confidence level typically declines gradually. The tracker may incorrectly predict the position of the object because of the confident error, causing a drift of the prediction box. To address this issue, when the box offset exceeds the threshold , we calculate the moving average of DC within the range of the past time sliding window [51] to capture the trend of confidence. A gradual decrease in DC indicates that confident errors will occur. We use the object state predicted by the Kalman filter as the tracking result to handle this situation. In this way, the problem of interference from similar objects can be addressed.
@@ V-A Experiment Settings
HOTC-2020 Dataset: HOTC-2020 was proposed at the first Hyperspectral Object Tracking Challenge in 2020. It consists of 75 HS video sequences, of which 40 are used for training, and 35 are used for testing. The training part contains a total of 12,839 annotated frames, while the testing part contains 16,574 annotated frames.
HOTC-2023 Dataset: HOTC-2023 was proposed at the 2023 Hyperspectral Object Tracking Challenge, which contains three sub-datasets: 15-band in the wavelength from 460 to 600 nm, 16-band in the wavelength from 600 to 850 nm, and 25-band in the wavelength from 665 to 960 nm sets. Each frame in the subset is labeled with a regression box.
2) Evaluation Metrics. We employed two widely used evaluation metrics: the area under the curve (AUC) of the success plot and the precision rate at the threshold of 20 pixels to conduct experiments. These metrics are extensively utilized in the field to assess performance in challenging scenarios, including occlusions, and provide a robust measure of tracker accuracy and robustness. The success plots depict the percentage of frames that meet or exceed different IOU thresholds, forming a success curve with thresholds ranging from 0 to 1. The precision rate analyzes the pixel difference between the centers of the predicted and actual bounding boxes. The precision rate curve is generated by displaying the percentage of frames where the center discrepancy is less than a specified pixel threshold, which ranges up to 50 pixels. Typically, the precision plot with the threshold of 20 is used to rank the trackers, denoted as DP_20.
3) Implementation Details. The proposed SPDAN is implemented in PyTorch. All experiments are conducted on a system equipped with NVIDIA RTX 4090 GPU, Intel I7-13700K CPU, and 32G RAM. The backbone network SPBN and the embedding network of RGB modality are initialized with the pre-trained parameters in [46]. The AdamW optimizer with a weight decay of 0.0001 is adopted to optimize the parameters of the network. The template and search patches are set to pixels and pixels, respectively. The network is then trained for a total of 10 epochs. and the batch size is set to 32. Meanwhile, the thresholds for DC and offset are set to 0.1 and 20, respectively.
@@
| B+D | 0.697 | 0.929 | +0.004 | +0.005 |
| B+DA | 0.702 | 0.935 | +0.009 | +0.011 |
| B+DA+S | 0.725 | 0.968 | +0.032 | +0.044 |
V-B Ablation Experiments
1) Effectiveness of the Proposed SEN: The spectral embedding network is proposed to refine the spectral features and form a robust prompt representation. To demonstrate the effectiveness of the SEN, we first construct a prompt-based framework following [20], which is denoted as “B”. Then, the SEN is utilized to incorporate the encoded spectral information into the visual features during the token embedding stage. It is denoted as “S”. We conduct the comparative experiments between the two models on the BihoT and HOTC-2020 datasets. The experimental results are listed in Tables III and IV. It can be observed that the SEN significantly improves the performance compared to the baseline model. Specifically, on the BihoT dataset, the AUC score increases by 3.3%, and the DP_20 improves by 3.8%. On the HOTC-2020 dataset, the AUC score increases by 1.1%, and the DP_20 improves by 1.4%. This indicates that spectral information is important for the model to capture camouflaged objects,
2) Effectiveness of the Proposed DAM: DAM is proposed to perceive the occlusion distractors and rectify the tracking results. We add the DAM to the model to show its effectiveness. Specifically, the methods in [21] are denoted as “D”. The DAM is denoted as “DA”. The experimental results are listed in Tables III and IV. It can be observed that DAM further enhances the tracking performance. Specifically, it increases the AUC and DP_20 on the BihoT dataset from 0.639 and 0.767 to 0.670 and 0.799, respectively. Compared to the baseline model, there are improvements of 3.1% and 3.2% in both indicators. For the HOTC-2020 dataset, DAM increases the AUC score and DP_20 from 0.693 and 0.924 to 0.702 and 0.935, respectively. Compared with the baseline model, there are improvements of 0.9% and 1.1% in both indicators. These results validate the effectiveness of the proposed DAM, which can correct the tracking results misguided by the distractors through a proposed decision confidence metric and three-step processes. Furthermore, to investigate the influence of the DC threshold , we conduct experiments on the HOTC-2020 dataset. The experimental results are listed in Table V. We set multiple thresholds, including 0.05, 0.10, 0.15, and 0.20. It can be seen that the optimal parameter is 0.1. Under this setting, the tracker can perceive distractors accurately, effectively improving the performance of the tracker.
| AUC | DP_20 | (AUC) | (DP_20) | |
| 0.05 | 0.712 | 0.950 | - | - |
| 0.10 | 0.725 | 0.968 | +0.013 | +0.018 |
| 0.15 | 0.701 | 0.934 | -0.011 | -0.016 |
| 0.20 | 0.695 | 0.923 | -0.017 | -0.027 |
@@ V-B Ablation Experiments
| Test setting | AUC | DP_20 | (AUC) | (DP_20) |
| B to B | 0.672 | 0.805 | - | - |
| N to B | 0.322 | 0.394 | -0.350 | -0.411 |
| N to N | 0.709 | 0.901 | - | - |
| B to N | 0.619 | 0.829 | -0.090 | -0.072 |
3) The Impact of Embedding Dimension In the SEN: To explore the impact of embedding dimension (the number of output channels by 3-D convolutional layer) on the performance of the proposed SEN, we select multiple embedding dimension settings and conduct the tracking performance experiments on the proposed BihoT. The experimental results are listed in Table VI. It can be seen that the four sets of embedding parameters (i.e., 2, 4, 8, and 16) are selected. Among them, the setting of embedding dimension 16 achieves the best results: the AUC score reaches 0.672, and DP_20 reaches 0.805. Under optimal experimental settings, the SEN can capture more comprehensive discriminative information in the spectral dimension, providing richer spectral semantic information for the prompt module.
4) The Generalization Ability Validation of SPDAN. To demonstrate the robustness of our method and the suitability of the dataset size, we perform cross-dataset validation. Specifically, our experiments involve two datasets, BihoT and HOTC23-NIR, each consisting of 25 bands. Our method entails training the SPDAN on the HOTC23-NIR training dataset and evaluating its performance on the BihoT testing dataset. Conversely, we also train the SPDAN on the BihoT training dataset and assess its performance on the HOTC23-NIR testing dataset. The experimental results are shown in Table VII and Fig. 7. The results indicate that the model trained on the BihoT dataset performs better on the HOTC23-NIR testing dataset. This outcome serves as evidence of the generalization capability of the SPDAN model and validates the adequacy of the dataset scale for training models with broad applicability. Moreover, an interesting phenomenon is observed: the model trained on HOTC23-NIR struggles to track many camouflaged objects in the BihoT testing set. This suggests that models trained on the existing HOT dataset face challenges in handling camouflaged objects and effectively utilizing HS information to meet practical application needs, such as tracking military camouflaged objects. This underscores the necessity and feasibility of the dataset we propose.
| Trackers | OSTrack[46] | SiamBAN[52] | MixFormer[38] | SEE-Net[16] | SiamHYPER[14] | SAM-T [53] | SPDAN (Ours) |
| AUC | 0.350 | 0.343 | 0.388 | 0.315 | 0.403 | 0.478 | 0.677 |
| DP_20 | 0.438 | 0.445 | 0.454 | 0.422 | 0.539 | 0.589 | 0.811 |
@@ V-C Comparison with State-of-the-arts
| Band Num. | Metrics | SPIRIT | SENSE | BAE-Net | SiamBAG | SEE-Net | SiamHYPER | SiamCAT | SPDAN (Ours) |
| 15 | AUC | 0.381 | 0.394 | 0.397 | 0.374 | 0.363 | 0.430 | 0.480 | 0.554 |
| DP_20 | 0.501 | 0.500 | 0.508 | 0.517 | 0.496 | 0.579 | 0.665 | 0.694 | |
| 16 | AUC | 0.608 | 0.607 | 0.558 | 0.554 | 0.561 | 0.590 | 0.593 | 0.669 |
| DP_20 | 0.820 | 0.826 | 0.802 | 0.779 | 0.784 | 0.815 | 0.822 | 0.888 | |
| 25 | AUC | 0.623 | 0.545 | 0.438 | 0.485 | 0.455 | 0.515 | 0.665 | 0.714 |
| DP_20 | 0.838 | 0.771 | 0.779 | 0.751 | 0.753 | 0.852 | 0.920 | 0.923 |
@@
Performance on the BihoT Dataset. We compare the proposed SPDAN with advanced HS and RGB trackers on the BihoT dataset, including SEE-Net [16], SiamHYPER [14], MixFormer [38], SiamBAN [52], SAM-T [53] and OSTrack [46]. SEE-Net reconstructs the HS bands through an autoencoder to learn the weight of each band. SiamHYPER proposes a Siamese tracking network that adopts independent networks for the extraction of features of RGB and HS modalities. MixFormer proposes a mixed attention module to extract features, and the tracking box can be obtained without a certain head network. SiamBAN proposes a box-adaptive head network to predict the accurate object positions. SAM-T amalgamates Segment Anything Model (SAM) with a proposed AOT-based tracking model to facilitate object tracking in video, which is built upon large visual models. OSTrack is a one-stage tracking framework, which extracts the features of the template and search images simultaneously through a visual Transformer-based backbone. It is worth noting that we retrained the SEE-Net and SiamHYPER according to the corresponding settings in the paper. The experimental results are listed in Table VIII. It can be observed that the proposed SPDAN achieves the best performance in both metrics. Furthermore, the BihoT dataset contains a large number of camouflage objects, which exhibit the same appearances but different spectral information. The lack of visual features poses a significant challenge to traditional RGB trackers that rely on visual features such as color and texture. This unique setting causes substantial disparities in AUC score and DP_20 between the conventional RGB trackers and the proposed SPDAN. This discrepancy not only highlights the limitations of current trackers but also emphasizes the effectiveness of our methods. In addition, most HS trackers use band regrouping to generate three-channel false-color images and directly use RGB trackers to predict object positions without considering spectral information. They heavily rely on the performance of RGB trackers and are not conducive to improving the performance of HS tracking. However, SiamHYPER achieves commendable results. Unlike band regrouping-based methods, it extracts spectral features and fuses cross-modality features, which, to some extent, utilizes spectral information and alleviates the dependence on RGB trackers. It indicates the importance of spectral information for HOT, especially for camouflaged objects.
@@
Performance on the HOTC-2020 Dataset. In this sub-section, we compare the proposed SPDAN with recent HS and RGB trackers in this sub-section. HS trackers include BAE-Net [59], SST-Net [60], SiamHYPER [14], SEE-Net [16], SiamOHOT [15], SPIRIT[13], MMF-Net [28], and SENSE [12], which are tested on HS video sequences. The RGB trackers include SiamBAN [52], SiamGAT [54], SiamCAR [55], SimTrack [56], OSTrack [46], SeqTrack [57], ARTrack [58], and MixFormer [38], which are tested on false-color video sequences. The experimental results are shown in Table IX. OSTrack* represents conducting experiments using weights trained in 300 epochs, while OSTrack represents conducting experiments using weights trained in 100 epochs. The proposed SPDAN can achieve results of 0.725 on AUC and 0.968 on DP_20, surpassing the above recent methods and proving the effectiveness of our method. The AUC of RGB trackers is mostly below 0.600, and the DP_20 is below 0.900. The performance of SPDAN is superior to the methods with the same ViT backbone, proving the effectiveness of modality fusion and DAM in improving HOT performance. The performance of HS trackers is generally better than that of RGB trackers. This result proves the necessity of leveraging spectral features, as relying solely on visual features cannot handle distractors from similar objects and background clutter. Then, comparing with HS trackers, it can be observed that the use of cross-modality feature fusion performs better, indicating that fusing RGB and HS modalities can help improve performance. In addition, most existing HS trackers are based on the Siamese architecture of SiamFC [61], ignoring the advantages of the Transformer in the feature extraction. Their global relationship modeling ability is very helpful for extracting the spatial and spectral information. However, ViT-based trackers like SimTrack and SeqTrack face limitations due to their heavy reliance on visual features and a semantic discrepancy with the robust HS features, impeding their performance enhancement. Their AUC yields 0.600 and 0.594, while the DP_20 are 0.845 and 0.856, respectively. In comparison, our SPDAN has achieved improvements of 12.5% and 13.1% in AUC and 12.3% and 11.2% in DP_20, respectively.
Performance on HOTC-2023 Dataset. Furthermore, to validate the effectiveness of our method, we conduct a comparative analysis of the proposed SPADN with several other trackers, including SiamCAT [27], SPIRIT [13], SENSE [12], BAE-Net [59], SiamBAG [26], and SiamHYPER [14], all of which are evaluated on HS images. The experimental results are listed in Table X. It is evident from the findings that our SPDAN model outperforms all other trackers in tracking performance across all three bands of HS data. Notably, trackers such as BAE-Net, SiamBAG, and SEE-Net rely on band regrouping techniques using false-color images, which can result in varying image quality across bands and consequently impact tracking accuracy. In contrast, methods like SiamHYPER and our proposed SPDAN, which focus on cross-modality feature fusion, demonstrate superior results, underscoring the effectiveness of modality fusion in enhancing feature robustness.
@@
To demonstrate the superiority of our method, we show the tracking results on the proposed BihoT dataset. We select four video sequences, including lemon2, apple3, chill2, and tank1. Due to the lack of RGB images collected in our dataset, we select the 1st, 9th, and 16th bands to form the false-color image. The experimental results are shown in Fig. 8. The ground truth bounding box is framed by black lines. The first column is obtained from the second frame of tracking results. The results from the consecutive frames are illustrated from the second to the fourth columns. From the first and second video sequences, we can observe that most methods struggle to accurately distinguish between camouflaged and real objects, resulting in tracking failures. In contrast, our method successfully captures both visual and spectral information of the objects, demonstrating the robustness in handling camouflaged objects. In addition, for the third video sequence, SPDAN can still perform well for smaller objects, especially when dealing with object rotation. Furthermore, the fourth sequence introduces two types of challenges related to camouflage: the first is the environmental similarity to the tank model, and the second is the visual similarity between two tanks made from different materials. It can be seen that the proposed method can effectively address these two challenges. In comparison, methods like SiamHYPER and MixFormer are misled by the camouflaged tanks during object movement. Other methods are also impacted by background clutter or camouflaged elements, further underscoring the superior performance of our method. In addition, we can also see typical failure cases and challenges, as shown in Figure 9. Due to the manual control of the object’s movement, the annotation box may contain some information about the attached objects, such as hands. When disturbed by occlusion or other factors, the tracker may focus on other areas of the hand. How to extract more refined features and avoid this bias is a major challenge for future research.
V-E Running Cost Comparison.
To demonstrate the superiority of our method in terms of running speed, we conduct experiments on the BihoT dataset. The experimental results using frames per second (FPS) as the evaluation indicator are reported in Table XI. All the trackers ran on a Windows machine with the configuration of an Intel(R) Core(TM) i7-13700K CPU@3.40GHz and an NVIDIA 4090 GPU. Besides, the RGB trackers ran on the three-channel false-color images, and the HS tracker ran on the HSIs. SEE-Net, SiamHYPER, and SPDAN are HS trackers, and others are RGB trackers. SPDAN achieves the best running speed among all trackers. It proves that our method has better computational efficiency compared to band regrouping or existing modality fusion methods. In addition, we also conducted experiments on computational complexity, and the experimental results are listed in Table XII. The method based on band regrouping requires multiple inferences and has higher flops. Meanwhile, the proposed SPDAN, utilizing 3D convolution, does not achieve the lowest FLOPs or model size but represents a balanced trade-off for performance improvement.